CARBON MOLECULE CHEMISTRY: GENERATION AND SPECTROSCOPIC CHARACTERIZATION OF TRIPLET HC_{5}H AND SUBSTITUTED ANALOGUES

\underline{PHILLIP S. THOMAS}, NATHAN P. BOWLING, and ROBERT J. McMAHON, Department of Chemistry, University of Wisconsin, Madison, WI 53706.

Triplet HC_{5}H,^{a} along with several derivatives (RC_{5}R′; R, R′ = H-, CH_{3}-, (CH_{3})_{3}C-), have been observed and characterized by IR, UV-vis, and ESR spectroscopy in cryogenic matrices. The structural similarity of RC_{5}R molecules to known interstellar molecules allows their consideration as plausible interstellar species. The electronic spectrum of HC_{5}H clarifies some ambiguities concerning earlier attempts to detect this species and provides a foundation for further gas-phase spectroscopic studies. The optical spectra of HC_{5}H and analogs also exhibit rich vibronic fine structure, and will be discussed in the context of the diffuse interstellar bands (DIBs). In addition, the photoisomerizations of RC_{5}R molecules will be briefly discussed. aN. P. Bowling, R. J. Halter, J. A. Hodges, R. A. Seburg, P. S. Thomas, C. S. Simmons, J. F. Stanton, and R. J. McMahon. J. Am. Chem. Soc. (2006) ASAP.